1. In the laboratory you dilute 3.81 mL of a concentrated 6.00 M perchloric acid solution to a total volume of 175 mL. What is the concentration of the dilute solution? ___M

2. In the laboratory you dilute 3.15 mL of a concentrated 3.00 M hydrobromic acid solution to a total volume of 100 mL. What is the concentration of the dilute solution? ___M
1
Expert's answer
2019-10-10T06:31:51-0400

The Molarity of new solution can be calculated using the formulae :

M1V1 = M2V2

where M1 = Initial molarity, M2= final molarity , V1= Initial volume, V2= Final volume

(1) M1= 6 M, M2= ? , V1= 3.81 mL, V2= 175 mL

Thus : 6 * 3.81 = M2 * 175

M2 =0.1306 M


(2) M1= 3 M, M2= ? , V1= 3.15 mL, V2= 100 mL

Thus : 3 * 3.15 = M2 * 100

M2 =0.0945 M